OpenStudy (b2st):

I'm stuck! Need help with rationalizing... Problem: if x = root 6 and y^2 =12, then (4/xy) = This is what I have so far... I made y = 2 root 6 so it correspond x= root 6 then I plugged in the x and y. However, I'm stuck at rationalizing it. (4)/ (2 root 6 * 2 root 6) = Idk if I'm supposed to do this... (4) / (4 root 6) I know I have to rationalize it something with root 6, but I don't get the same result in all the multiple choice.. I probably dome something wrong.. So if you can please tell me step by step it would be nice ^ ^

OpenStudy (b2st):

I know I have to rationalize it with root 6, by multiplying both root 6 top and bottom...

OpenStudy (yttrium):

Do you mean x = square root of 6??

OpenStudy (b2st):

yup. oops. disregard the y = 2 square root 6.....

OpenStudy (b2st):

lol. scrap everything I said, but I'm still stuck on this problem...

OpenStudy (yttrium):

Well try to look at this. Problem: if x = root 6 and y^2 =12, then (4/xy) 4/xy = 4/[(sqrt6)(sqrt12) = 4/sqrt(72) = 4/sqrt (9 * 2^3) = 4/[2(3sqrt(2))] = 2/3sqrt(2) -> rationalizing by multiplying num and denum by sqrt(2) = 2sqrt(2)/6 Therefore, 4/xy = sqrt(2)/3

OpenStudy (b2st):

Thank youuu! I was over thinking this problem too much. whoops.

OpenStudy (yttrium):

No problem. Just post again, if you encounter new confusion. :D

OpenStudy (b2st):

Oh, I'm confuse about = 4/sqrt (9 * 2^3) = 4/[2(3sqrt(2))] How and why did this happen?

OpenStudy (b2st):

wait never mind.. lol